Find the equation for the line that passes through the point (-2,4), and that is parallel to the line with the equation y=3/4×-7/2
Answer by stanbon(75887)

Find the equation for the line that passes through the point (-2,4), and that is parallel to the line with the equation y=3/4×-7/2
------
slope = 3/4
---
Form: y = mx + b
Solve for "b":
4 = (3/4)(-2) + b
b = 4+(3/2) = 11/2
----
Equation:
y = (3/4)x + 11/2
===================
Cheers,
Stan H.
===================